OpenStudy (anonymous):

What are the zeros of x^6-64 ?

OpenStudy (anonymous):

The zeros are the x-interceps

OpenStudy (anonymous):

intercepts

OpenStudy (anonymous):

You can use the quadratic formula to find the zeros.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

OK, so the quadratic formula is -b+- sqrt(b^2-4ac) all over 2a

OpenStudy (anonymous):

(8 + x3)(-8 + x3)

hero (hero):

Simply using the formula for difference of squares would have been sufficient enough. \(a^2 + b^2 = (a + b)(a - b)\) \(x^6 - 64 = (x^3)^2 - 8^2 = (x^3 + 8)(x^3 - 8)\)